# Team Westwood Solution for MIDOG 2025 Challenge: An Ensemble-CNN-Based Approach For Mitosis Detection And Classification

### v2

@@ Abstract
This abstract presents our solution (Team Westwood) for mitosis detection and atypical mitosis classification in the MItosis DOmain Generalization (MIDOG) 2025 challenge (1). For mitosis detection, we trained an nnUNetV2 for initial mitosis candidate screening with high sensitivity, followed by a random forest classifier ensembling predictions of three convolutional neural networks (CNNs): EfficientNet-b3, EfficientNet-b5, and EfficientNetV2-s. For the atypical mitosis classification, we trained another random forest classifier ensembling the predictions of three CNNs: EfficientNet-b3, EfficientNet-b5, and InceptionV3. On the preliminary test set, our solution achieved an F1 score of 0.7450 for track 1 mitosis detection, and a balanced accuracy of 0.8722 for track 2 atypical mitosis classification. On the final test set, our solution achieved an F1 score of 0.6972 for track 1 mitosis detection, and a balanced accuracy of 0.8242 for track 2 atypical mitosis classification.
{corrauthor}
hgu2@kumc.edu
Introduction
@@
In pathology, mitosis activity assessment in the Hematoxylin and Eosin (H&E) slides by human pathologists can be challenging due to its small size and low prevalence in low-grade tumors (2, 3). Recent advancements in digital pathology and artificial intelligence (AI) can provide a low-cost computer-assisted solution for more timely and precise examination (4, 5). Despite this, perhaps one hurdle for AI applicability is its generalizability on high variance of pathology datasets, due to three factors: (1) the intrinsic appearance difference of mitosis and their mimickers across tumor types; (2) processing protocols from different labs; and (3) scanner imaging settings and image post-processing algorithms.
@@
To fill this gap, several large-scale mitosis datasets covering various organs, scanners, and atypical mitotic figures have been recently curated and made publicly available (6, 7, 8, 9). Therefore, re-training AI models on these new datasets and running a more comprehensive evaluation has become increasingly necessary. In MIDOG 2022, we employed an EfficientNet-b3 CNN for both detection and classification of mitosis (10). While this design was compact in terms of model parameters, it relied on calculating attentions for mitosis localization, which could not be easily parallelized and thus had limited efficiency.
@@
As an improvement, in MIDOG 2025 (1), we adopted the latest nnUNetV2111https://github.com/MIC-DKFZ/nnUNet/tree/master/nnunetv2 as a lightweight and fast mitosis candidate localization and screening module. During training, both true positives and hard negatives were treated as positive samples to enhance its sensitivity. For each mitosis candidate, we then applied a “heavier” random forest of three CNN models (i.e., EfficientNet-b3, EfficientNet-b5, and EfficientNetV2-s) to achieve specificity. For track 2 atypical mitosis classification task, we also used a random forest ensembling EfficientNet-b3, EfficientNet-b5, and InceptionV3, aiming to achieve more robust performance.
Methods
@@ 0.1 Track 1: Mitosis detection
AI Pipeline Following the popular solutions in MIDOG 2022 (11), we designed a two-stage mitosis segmentation – verification pipeline to balance inferencing efficiency and detection performance. Specifically, we used the nnUNetV2 for stage-1 segmentation and a random forest of three CNNs for stage-2 verification, as shown in Figure 1.
Dataset We included MIGOG++ (8), MITOS_WSI_CMC (7), and MITOS_CCMCT (6) for model training and validation (70,724 mitoses in total). Approximately $\sim$90% of the slides or regions of interest (ROIs) were used for model training, and the rest for validation. To train the nnUNetV2, we randomly cropped 253,703 ($\times512$-pixel) patches from the training slides/ROIs. Both ground-truth mitoses and hard-negative mimickers were treated as positives (to improve the sensitivity, for nnUNetV2 training only). For each positive, we synthesized the segmentation mask by drawing a filled circle (45-pixel radius) centered at its location. The trained nnUNetV2 with the best sensitivity was then applied to both training and validation slides/ROIs. From all segmentation hotspot centroids, we extracted $\times140$-pixel patches (141,224 positives and 2,044,045 negatives [^2] The positives consist of 70,971 samples from ground-truth, 67,206 true positive predictions, and 3,047 false negatives by nnUNetV2. The negatives are false positive predictions by nnUNetV2.) for subsequent CNN training.
@@
Model Training and Validation Firstly, the segmentation model (nnUNetV2) was trained with oversample foreground percent 50%, initial learning rate 0.001, weight decay $10^{-4}$, AdamW optimizer, DICE loss, Cosine Annealing LR with Warm Restarts (T_0: 10, T_mult: 1) for 50 epochs. Data augmentation included random image transform (e.g., crop, scaling, rotation, flip, mirror), color intensity transform (e.g., brightness, contrast, and gamma adjustments), random gaussian noise and gaussian blur. After each training epoch, the checkpoint was evaluated on the entire validation set of slides/ROIs, and sensitivity was calculated. The checkpoint with the highest sensitivity was selected for final inferencing.
For the classification models (multiple CNNs) training, we used the initial learning rate $8\times 10^{-4}$, weight decay $10^{-4}$, AdamW optimizer, cross entropy loss, Cosine Annealing LR with Warm Restarts (T_0: 15, T_mult: 1) for 80 epochs. Data augmentation includes random image transform (e.g., crop, flip, rotation), color adjustment (e.g., brightness, hue, saturation), random gaussian noise and gaussian blur. We tried eight CNN variants: EfficientNet-b3, EfficientNet-b5, EfficientNetV2-s, EfficientNetV2-m, InceptionV3, ResNeXt50_32x4d, ViT-b, and SwinV2-s. After each epoch, the checkpoint of each CNN was evaluated on the extracted validation patches. We trained each CNN for 80 epochs, and the top three CNNs (i.e., EfficientNet-b3, EfficientNet-b5, and EfficientNetV2-s) with the highest F1 scores were selected to construct the final ensemble.
Inferencing and Ensembling Training Test-time augmentation (TTA) was applied to both nnUNetV2 ($\times$3; random flip and rotation) and each of the three CNNs ($\times$3; central random crop, flip, and rotation). For each candidate prediction, this TTA generated nine probability outputs (3 CNNs $\times$ 3 TTA). A random forest classifier (n_estimators=260, max_depth=1) was then trained using the output probabilities from all TTA results for each candidate as input features to estimate the final prediction probability. For submission, the pipeline was run on the test images using a 512-pixel sliding-window with 256-pixel overlap.
@@ 0.2 Track2: Atypical mitosis classification
AI Pipeline A random forest of three CNNs (see Figure 2) was used to improve performance due to the small training set.
Dataset AMi-Br (12) and MIDOG 2025 Atypical Training Set (13) (13,077 mitoses and 2,580 atypical mitoses) were included. Approximately 85% of the dataset was used for model training, and the rest for validation and threshold selection. All images were rescaled to $\times128$ pixels for training and inferencing.
Model Training and Validation Similar to track 1-CNN, we trained eight CNN variants with the same hyperparameters and data augmentation strategy: EfficientNet-b3, EfficientNet-b5, EfficientNetV2-s, EfficientNetV2-m, Inception_V3, ResNeXt50_32x4d, ViT-b, and SwinV2-m. Three CNNs of EfficientNet-b3, EfficientNet-b5, and Inception_V3 were selected because they achieved the highest balanced accuracies during validation.
Inferencing and Ensembling Training For each CNN, TTA ($\times$5; random flip and rotation) was used during the inferencing. An ensemble module made the final prediction by averaging the 15 concatenated probability outputs with equal weights, as this approach outperformed the random forest method in the Track 2 preliminary test.
Results
@@
Discussion
The results demonstrate that an ensemble of CNN classifiers can contribute to the AI performance when combined with a lightweight segmentation model and when used independently for classification, where random forest is a good option to support the ensemble module. We notice that during the validation step of both tracks, the F1 score becomes less sensitive to the threshold with the ensemble module, as it exhibits a wider range of plateau in the F1-threshold plot, demonstrating the stability of our approaches as the most significant advantage. This is further proven by the fact that our algorithm for track 1 is ranked at #19 in preliminary test (included <100 hotspot ROIs) but is then ranked at #7 in final test (with random selections from the WSIs).
However, the relatively low F1 scores and recall for certain tumor types in both preliminary test and final test indicate that challenges still remain in generalization. Future work may focus on enhancing hard example mining, such as incorporating additional mitosis-like negative samples, to further improve the model’s ability to distinguish difficult cases.

### v3

@@ Abstract
This abstract presents our solution (Team Westwood) for mitosis detection and atypical mitosis classification in the MItosis DOmain Generalization (MIDOG) 2025 challenge [1]. For mitosis detection, we trained an nnUNetV2 for initial mitosis candidate screening with high sensitivity, followed by a random forest classifier ensembling predictions of three convolutional neural networks (CNNs): EfficientNet-b3, EfficientNet-b5, and EfficientNetV2-s. For the atypical mitosis classification, we trained another random forest classifier ensembling the predictions of three CNNs: EfficientNet-b3, EfficientNet-b5, and InceptionV3. On the preliminary test set, our solution achieved an F1 score of 0.7450 for track 1 mitosis detection, and a balanced accuracy of 0.8722 for track 2 atypical mitosis classification. On the final test set, our solution achieved an F1 score of 0.6972 for track 1 mitosis detection, and a balanced accuracy of 0.8242 for track 2 atypical mitosis classification.
1 Introduction
@@
In pathology, mitosis activity assessment in the Hematoxylin and Eosin (H&E) slides by human pathologists can be challenging due to its small size and low prevalence in low-grade tumors [13, 6]. Recent advancements in digital pathology and artificial intelligence (AI) can provide a low-cost computer-assisted solution for more timely and precise examination [5, 11]. Despite this, perhaps one hurdle for AI applicability is its generalizability on high variance of pathology datasets, due to three factors: (1) the intrinsic appearance difference of mitosis and their mimickers across tumor types; (2) processing protocols from different labs; and (3) scanner imaging settings and image post-processing algorithms.
@@
To fill this gap, several large-scale mitosis datasets covering various organs, scanners, and atypical mitotic figures have been recently curated and made publicly available [7, 2, 4, 14]. Therefore, re-training AI models on these new datasets and running a more comprehensive evaluation has become increasingly necessary. In MIDOG 2022, we employed an EfficientNet-b3 CNN for both detection and classification of mitosis [9]. While this design was compact in terms of model parameters, it relied on calculating attentions for mitosis localization, which could not be easily parallelized and thus had limited efficiency.
@@
As an improvement, in MIDOG 2025 [1], we adopted the latest nnUNetV2111https://github.com/MIC-DKFZ/nnUNet/tree/master/nnunetv2[12] as a lightweight and fast mitosis candidate localization and screening module. During training, both true positives and hard negatives were treated as positive samples to enhance its sensitivity. For each mitosis candidate, we then applied a “heavier” random forest of three CNN models (i.e. EfficientNet-b3, EfficientNet-b5, and EfficientNetV2-s) to achieve specificity. For track 2 atypical mitosis classification task, we also used a random forest ensembling EfficientNet-b3, EfficientNet-b5, and InceptionV3, aiming to achieve more robust performance.
2 Methods
@@ 2.1 Track 1: Mitosis detection
2.1.1 AI Pipeline
Following the popular solutions in MIDOG 2022 [3], we designed a two-stage mitosis segmentation– verification pipeline to balance inferencing efficiency and detection performance. Specifically, we used the nnUNetV2 for stage-1 segmentation and a random forest of three CNNs for stage-2 verification, as shown in Figure 1.
2.1.2 Dataset
We included MIDOG++[4], MITOS_WSI_CMC[2], and MITOS_C CMCT[7] for model training and validation (70,724 mitoses in total). Approximately $\sim$90% of the slides or regions of interest (ROIs) were used for model training, and the rest for validation. To train the nnUNetV2, we cropped 253,703 ($\times512$-pixel) patches from the training slides/ROIs, including positive patches randomly cropped around ground truth labels and negative patches randomly cropped from background. Both ground-truth mitoses and hard-negative mimickers were treated as positives (to improve the sensitivity, for nnUNetV2 training only). For each positive, we synthesized the segmentation mask by drawing a filled circle (45-pixel radius) centered at its location. The trained nnUNetV2 with the best sensitivity was then applied to both training and validation slides/ROIs. From all segmentation hotspot centroids, we extracted $\times140$-pixel patches (141,224 positives and 2,044,045 negatives [^2] The positives consist of 70,971 samples from ground-truth, 67,206 true positive predictions, and 3,047 false negatives by nnUNetV2. The negatives are false positive predictions by nnUNetV2.) for subsequent CNN training.
@@ 2.1.2 Dataset
2.1.3 Model Training and Validation
Firstly, the segmentation model (nnUNetV2) was trained with oversample foreground percent 50%, initial learning rate 0.001, weight decay $10^{-4}$, AdamW optimizer, DICE loss, Cosine Annealing LR with Warm Restarts (T_0: 10, T_mult: 1) for 50 epochs. Data augmentation included random image transform (e.g. crop, scaling, rotation, flip, mirror), color intensity transform (e.g. brightness, contrast, and gamma adjustments), random gaussian noise and gaussian blur. After each training epoch, the checkpoint was evaluated on the entire validation set of slides/ROIs, and sensitivity was calculated. The checkpoint with the highest sensitivity was selected for final inferencing.
For the classification models (multiple CNNs) training, we used the initial learning rate $8\times 10^{-4}$, weight decay $10^{-4}$, AdamW optimizer, cross entropy loss, Cosine Annealing LR with Warm Restarts (T_0: 15, T_mult: 1) for 80 epochs. Data augmentation includes random image transform (e.g. crop, flip, rotation), color adjustment (e.g. brightness, hue, saturation), random gaussian noise and gaussian blur. We tried eight CNN variants: EfficientNet-b3, EfficientNet-b5, EfficientNetV2-s, EfficientNetV2-m, InceptionV3, ResNeXt50_32x4d, ViT-b, and SwinV2-s. After each epoch, the checkpoint of each CNN was evaluated on the extracted validation patches. We trained each CNN for 80 epochs, and the top three CNNs (i.e. EfficientNet-b3, EfficientNet-b5, and EfficientNetV2-s) with the highest F1 scores were selected to construct the final ensemble.
2.1.4 Inferencing and Ensembling Training
Test-time augmentation (TTA) was applied to both nnUNetV2 ($\times$3; random flip and rotation) and each of the three CNNs ($\times$3; central random crop, flip, and rotation). For each candidate prediction, this TTA generated nine probability outputs (3 CNNs $\times$ 3 TTA). A random forest classifier (n_estimators=260, max_depth=1) was then trained using the output probabilities from all TTA results for each candidate as input features to estimate the final prediction probability. For submission, the pipeline was run on the test images using a 512-pixel sliding-window with 256-pixel overlap.
@@ 2.2 Track2: Atypical mitosis classification
2.2.1 AI Pipeline
A random forest ensembe of three CNNs (see Figure 2) was used to improve performance due to the small training set.
2.2.2 Dataset
AMi-Br [8] and MIDOG 2025 Atypical Training Set [15] (13,077 mitoses and 2,580 atypical mitoses) were included. Approximately 85% of the dataset was used for model training, and the rest for validation and threshold selection. All images were rescaled to $\times128$ pixels for training and inferencing.
2.2.3 Model Training and Validation
Similar to track 1-CNN, we trained eight CNN variants with the same hyperparameters and data augmentation strategy: EfficientNet-b3, EfficientNet-b5, EfficientNetV2-s, EfficientNetV2-m, Inception_V3, ResNeXt50_32x4d, ViT-b, and SwinV2-m. Three CNNs of EfficientNet-b3, EfficientNet-b5, and Inception_V3 were selected because they achieved the highest balanced accuracies during validation.
2.2.4 Inferencing and Ensembling Training
For each CNN, TTA ($\times$5; random flip and rotation) was used during the inferencing. An ensemble module made the final prediction by averaging the 15 concatenated probability outputs with equal weights, as this approach outperformed the random forest method in the Track 2 preliminary test.
3 Results
@@
4 Discussion
The results demonstrated that an ensembling module using random forest can improve the robustness and generalizability of AI detection. During our validation stage, we observed improvements in F1 score robustness when the number of trees (i.e., n_estimators) increased from 100 to 500. The result suggests that incorporating the ensembling approach can enable the mitosis detection pipeline to be less subject to the threshold selection across different organs, patients, and scanners during practical usage.
Based on the result, we suggest two directions for future improvements:
  1. 1.
    Data quality. It is noteworthy that all current public mitosis detection dataset was scanned using a single $z$ layer. In the meantime, Z-stacked scanning, which has become more available recently, can improve the whole slide imaging quality by preserving additional depth information in the $z$ direction. In our previous study [10], we observed that using z-stacked scans of five layers improved AI mitosis detection recall by 17.14%, while only having a marginal impact on the precision. Future work may explore similar techniques to improve the whole slide imaging quality (in both xyx-y and $z$ directions) and measure the corresponding effectiveness.
  2. 2.
    Data quantity and variation. During the final test phase of track 1, we noticed that the F1 score achieved by the 3rd solution (0.7085) to 10th (0.6883) only varied by $\sim 2\%$. Therefore, modifications on AI models or detection pipelines may not be the most effective way to further improve the performance: the focus may shift to extend the training dataset, such as incorporating additional hard negative samples, to further improve the model’s capability to distinguish the challenging, confusing patterns.